OpenStudy (anonymous):

write the equation in slope intercept form the line -7 and y intercept 6

myininaya (myininaya):

Is -7 the slope?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the slope is -7 and y intercept 6

myininaya (myininaya):

So do you know the slope intercept form?

OpenStudy (anonymous):

i sure dont the question was "write the equation in slope-intercept from the line that has a slope -7 and y intercept 6"

myininaya (myininaya):

y=mx+b is the slope-intercept form because m is the slope and b is the y-intercept

myininaya (myininaya):

So basically you were given m and b So just replace m and b and you got it! :)

OpenStudy (anonymous):

thank u

myininaya (myininaya):

np

OpenStudy (anonymous):

i just dont understand slopes =/

myininaya (myininaya):

Like how to use the slope in graphing?

OpenStudy (anonymous):

yeap.... its so confusing to me

myininaya (myininaya):

Slope=rise/run If the top number is negative, we go down If the top number is positive, we go up If the bottom number is negative, we go left If the bottom number is positive, we go right

OpenStudy (anonymous):

gosh you made it easy!! =)

myininaya (myininaya):

Coolness.

myininaya (myininaya):

So if we have y=6x-4 Where the y-intercept is -4 and the slope is 6 ( or 6/1 since any integer can be written as a fraction by doing that integer/1) We graph the y-intercept first (or called b by the y=mx+b form) Then use the slope to find another point on the graph So using the (0,-4) as a starting point, we find another point by going up 6 units from there and going over 1 unit since our slope was 6/1.

myininaya (myininaya):



myininaya (myininaya):